14. The combustion of a hydrocarbon produces 66.02 g CO2 (molecular weight 44.01 g/mol) and 13.51 g H2O (molecular weight 18.Please show all steps and explain, as I need to understand for an exam. Thx!
Answer #1

A hydrocarbon when burnt produces carbon dioxide and H2O

Given,mass of CO2 = 66.02 gm

Molar mass of of CO2 =44.01

Given mass of H2O = 13.51 gm

Molar mass of H2O =18.01

Number of moles of CO2 = given mass/molar mass =66.02/44.01 = 1.5 mol

Number of moles of H2O =13.51 /18.01 =0.75

In every reaction of combustion of hydrocarbon,production of CO2 and H2O is 1:2 ratio

Example--- CH4 + O2 ------> CO2 + 2H2O

Therefore above moles will be in 1:2 ratio

For 1.5 mol of carbon dioxide, (2× 0.75 ) 1.5mol of H2O is also produced

So ratio of number of moles of both are 1:1

Therefore empirical formula = CH